# Exotic Doublon-Holon Pairing State in Photodoped Mott Insulators

### v1

@@ Abstract
We demonstrate the existence of a unique pairing state in photodoped Mott insulators on ladder geometries, characterized by quasi-long-ranged doublon-holon correlations, using the density matrix renormalization group method. This phase exhibits doublon-holon pairing correlations with opposite signs along the rung and chain directions, reminiscent of $d$-wave pairing in chemically doped ladder systems. By constructing the phase diagram, we reveal that the doublon-holon pairing state emerges between the spin-singlet phase and the charge-density-wave/$\eta$-pairing phase. Our study suggests that the interplay of charge, spin, and $\eta$-spin degrees of freedom can give rise to exotic quantum many-body states in photodoped Mott insulators.
@@
External field driving is a useful tool for generating and controlling intriguing out-of-equilibrium phenomena in correlated electron systems [1, 2, 3, 4, 5, 6]. Photoinduced phase transitions [7, 8, 9, 10, 11, 12, 13], Floquet engineering [14, 15, 16, 17, 18], and high-harmonic generation [19, 20, 21, 22, 23, 24, 25, 26] have been extensively studied to date. In insulating materials, optical excitation across the band gap generates electron and hole carriers. When the recombination time is sufficiently long, the excited state potentially settles into a quasi-steady photodoped state, nearly conserving the number of carriers [27, 28]. Photoexcitation in correlation-driven Mott insulators (MIs) creates two types of carriers, doublons (doubly occupied sites) and holons (empty sites), see Fig. 1(b) [29, 30, 6, 31, 32, 33, 34], resulting in a unique photodoped state. In a large-gap MI, the carriers are expected to have long lifetimes, leading to a quasi-steady state where doublons and holons coexist [35, 36, 37, 38, 39, 40, 41, 42, 43, 44].
@@
Figure 1: (a) Phase diagram of the photodoped ladder-type MI as a function of the doublon density $n_{d}$ and the intersite Coulomb interaction $V$ with the blue, red, green, and purple points representing the predominance of the spin, $\eta$-pairing, CDW, and doublon-holon (DH) pairing correlations, respectively. At the orange point ($n_{d}=0.5$ and $V=0.2t_{\parallel}$), the $\eta$-pairing and CDW states are degenerate. Parameters: $t_{\perp}=t_{\parallel}$, $J_{\perp}=J_{\parallel}=0.4t_{\parallel}$, and $V=V_{\parallel}=V_{\perp}$. (b) DH pairing state in the photodoped ladder with doublons and holons generated by light. The DH pairing correlation is characterized by an operator that maps spin singlets on nearest-neighbor sites to DH $\eta$-triplets.
@@
Photodoping can be a trigger for creating exotic many-body states that are inaccessible in equilibrium, such as the $\eta$-pairing state [45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The photodoped states have been investigated in detail in the one-dimensional (1D) Hubbard chain, where the phase diagram comprises the spin density wave (SDW) phase at low doublon density $n_{d}$ and the charge density wave (CDW) and $\eta$-pairing phases from intermediate to high $n_{d}$ [43, 44]. Although even richer physical properties are anticipated in the two-dimensional (2D) systems because of the increased geometric degrees of freedom and the lack of spin-charge-$\eta$ separation [44], these properties of photodoped 2D MIs are yet to be understood.
@@
Ladder geometry can serve as a bridge between 1D and 2D systems and is more manageable for numerical approaches. For example, density matrix renormalization group (DMRG) calculations, which enable us to obtain numerically precise low-dimensional wave functions, have shown that chemically hole-doped ladders exhibit pair correlations corresponding to $d$-wave superconductivity (SC) in a 2D square lattice [57, 58, 59, 60, 61]. These studies are an important stepping stone to research on $d$-wave SC in the Hubbard and $t$-$J$ models on multileg cylinders, mimicking the 2D square lattice of cuprate superconductors [62, 63, 64, 65, 66, 67]. Furthermore, the pairing mechanism in ladder and bilayer systems [68, 69, 70, 71, 72, 73, 74, 75, 76, 77, 78, 79] has attracted significant attention associated with the recent discovery of high-temperature SC in bilayer nickelates [80, 81, 82, 83]. Because a strong interchain spin coupling favors the formation of spin-singlet rungs, hole doping to the ladder system leads to the formation of interchain hole-hole pairs to minimize the disruption of spin-singlet bonds, resulting in the development of SC correlations [68]. Realization of such a pairing mechanism has been recently proposed also with cold atoms in optical lattices [84]. From these considerations, we arrive at the crucial question: what are the equivalent paired states in photodoped MIs?
@@
In this Letter, we show that photodoped MIs in the ladder geometry and quasi-stationary situation exhibit an exotic pairing state composed of doublons and holons (Fig. 1). This pairing is characterized by the quasi-long-range correlation of doublon-holon (DH) pairs on the spin background shown in Fig. 1(b). We find opposite phases for the rung and chain pairing, analogous to signatures of $d$-wave pairing in chemically doped ladders. We reveal the roles of interchain coupling and propose an effective minimal model that can capture the pairing properties.
@@
Model and Methods—To model MIs in the presence of nonlocal interactions, we start with the extended Hubbard Hamiltonian in a ladder geometry
^Hub=subscript^$\hat{\mathcal{H}}_{\mathrm{Hub}}=$ $-t_{\parallel}\sum_{j,\alpha,\sigma}\left(\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j+1,\alpha;\sigma}+\mathrm{H.c.}\right)$
$-t_{\perp}\sum_{j,\sigma}\left(\hat{c}^{\dagger}_{j,0;\sigma}\hat{c}_{j,1;\sigma}+\mathrm{H.c.}\right)+U\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha;\uparrow}\hat{n}^{\prime}_{j,\alpha;\downarrow}$
$+V_{\parallel}\sum_{j,\alpha}\hat{n}^{\prime}_{j,\alpha}\hat{n}^{\prime}_{j+1,\alpha}+V_{\perp}\sum_{j}\hat{n}^{\prime}_{j,0}\hat{n}^{\prime}_{j,1}.$ (1)
@@
Here, $\hat{c}^{\dagger}_{j,\alpha;\sigma}$ ($\hat{c}_{j,\alpha;\sigma}$) is the creation (annihilation) operator for a fermion with spin $\sigma=\uparrow$, $\downarrow$ at site $j$ on chain $\alpha=0,1$. $\hat{n}_{j,\alpha;\sigma}=\hat{c}^{\dagger}_{j,\alpha;\sigma}\hat{c}_{j,\alpha;\sigma}$, $\hat{n}^{\prime}_{j,\alpha;\sigma}=\hat{n}_{j,\alpha;\sigma}-1/2$, and $\hat{n}^{\prime}_{j,\alpha}=\hat{n}^{\prime}_{j,\alpha;\uparrow}+\hat{n}^{\prime}_{j,\alpha;\downarrow}$. $t_{\parallel}$ and $t_{\perp}$ are the hopping integrals along the chain and rung directions, respectively. $U>0$ is the on-site Coulomb interaction. $V_{\parallel}>0$ and $V_{\perp}>0$ are the nearest-neighbor Coulomb interactions along the chain and rung directions, respectively.
In photodoped MIs, doublons and holons are generated by an optical excitation, as shown in Fig. 1(b). When $U\gg t_{\parallel},t_{\perp}$, the recombination time is sufficiently long [35, 36, 37, 38, 39, 41, 40, 42] and the system can be considered to be in a pseudoequilibrium state where doublons and holons are conserved [43, 44]. Assuming $U\gg t_{\parallel},t_{\perp}$, we introduce an effective model derived by the Schrieffer-Wolff transformation that excludes the hopping terms changing the number of doublons and holons [39, 41, 43, 85]. The Hamiltonian of the effective $t$-$J$-$V$ model is given by
^=^$\hat{\mathcal{H}}=\hat{\mathcal{H}}^{(0)}_{t}+\hat{\mathcal{H}}^{(s)}_{J}+\hat{\mathcal{H}}^{(\eta)}_{J}+\hat{\mathcal{H}}_{V}.$ (2)
Here, ^$\hat{\mathcal{H}}^{(0)}_{t}=\hat{\mathcal{H}}^{(0)}_{t_{\parallel}}+\hat{\mathcal{H}}^{(0)}_{t_{\perp}}$ denotes the doublon-number-conserving hopping, see Supplemental Material for details [86], and ^$\hat{\mathcal{H}}_{V}=\hat{\mathcal{H}}_{V_{\parallel}}+\hat{\mathcal{H}}_{V_{\perp}}$ represents the nearest-neighbor Coulomb interactions $V_{\parallel}$ and $V_{\perp}$ in Eq. (1). ^$\hat{\mathcal{H}}^{(s)}_{J}=\hat{\mathcal{H}}^{(s)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(s)}_{J_{\perp}}$ describes the Heisenberg-type spin-exchange interaction, while ^$\hat{\mathcal{H}}^{(\eta)}_{J}=\hat{\mathcal{H}}^{(\eta)}_{J_{\parallel}}+\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}$ represents the $\eta$-spin interaction for doublons and holons [43, 85]. These nearest-neighbor interactions between sites $\bm{i}$ and $\bm{j}$ are given by
^$\hat{\mathcal{H}}^{(s)}_{J_{\bm{i}\bm{j}}}=J_{\bm{i}\bm{j}}\left(\hat{\bm{s}}_{\bm{i}}\cdot\hat{\bm{s}}_{\bm{j}}-\frac{1}{4}\delta_{1,\hat{n}_{\bm{i}}\hat{n}_{\bm{j}}}\right),$ (3)
^$\hat{\mathcal{H}}^{(\eta)}_{J_{\bm{i}\bm{j}}}=-J_{\bm{i}\bm{j}}\left(\hat{\bm{\eta}}_{\bm{i}}\cdot\hat{\bm{\eta}}_{\bm{j}}-\frac{1}{4}\left(1-\delta_{1,\hat{n}_{\bm{i}}}\right)\left(1-\delta_{1,\hat{n}_{\bm{j}}}\right)\right),$ (4)
where $\bm{j}$ denotes $(j,\alpha)$. The spin operator is defined as $\hat{\bm{s}}_{\bm{j}}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{j,\alpha;\sigma}\bm{\sigma}_{\sigma,\sigma^{\prime}}\hat{c}_{j,\alpha;\sigma^{\prime}}/2$, where $\bm{\sigma}$ is the vector of Pauli matrices. The $\eta$-spin operator $\hat{\bm{\eta}}_{\bm{j}}$ is given by $\hat{\eta}^{+}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}^{\dagger}_{j,\alpha;\downarrow}\hat{c}^{\dagger}_{j,\alpha;\uparrow}$, $\hat{\eta}^{-}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$, and $\hat{\eta}^{z}_{\bm{j}}=\left(\hat{n}_{j,\alpha}-1\right)/2$ [87]. The explicit Hamiltonian for the ladder system is presented in the Supplemental Material [86]. For the hopping integral $t_{\bm{i}\bm{j}}$, the coupling constant is characterized by $J_{\bm{i}\bm{j}}=4t^{2}_{\bm{i}\bm{j}}/U>0$ (at $V_{\bm{i}\bm{j}}=0$). The $V$ term takes the form $V_{\bm{i}\bm{j}}\hat{n}^{\prime}_{\bm{i}}\hat{n}^{\prime}_{\bm{j}}=4V_{\bm{i}\bm{j}}\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}$, thereby introducing Ising-type anisotropy of $\eta$-spin. In contrast to the chemically doped $t$$J$ model, $\hat{\mathcal{H}}$ includes the interaction between doublons and holons described by $\eta$-spin, which enables the formation of $\eta$-pairing states in photodoped MIs [51, 43, 44].
@@
We describe the photodoped ladder as the lowest-energy state of $\hat{\mathcal{H}}$ with a fixed doublon number. We obtain the state using the DMRG method [88, 89, 90]. Unless otherwise specified, we use $t_{\parallel}$ as the unit of energy and set the chain length to $L=160$ with open boundary conditions. The total number of sites is $2L$, and the numbers of up- and down-spin fermions are set to $N_{\uparrow}=N_{\downarrow}=L$ (i.e., half-filling without spin polarization), while the state is configured as a function of the doublon number. The bond dimension is set to $m=10000$ with the maximum truncation error typically being on the order of $1.0\times 10^{-7}$.
Numerical Results—To investigate which nonequilibrium phases can emerge in photodoped ladders, we study the behavior of correlation functions $C(r)=\braket{\hat{O}^{\dagger}_{j_{0}+r}\hat{O}_{j_{0}}}$ between operators $\hat{O}_{j}$ that are $r$ sites apart. We calculate correlations with respect to the site $j_{0}=L/4+1$ to minimize open boundary effects. The dominant nonequilibrium phase is determined by identifying the correlation function that exhibits the slowest decay with spatial distance $r$. Because the correlations between the single-site operators are symmetric on both chains, we consider the $\alpha=0$ chain index in the operators. We use $\hat{O}_{j}=\hat{n}_{j,\alpha;\uparrow}-\hat{n}_{j,\alpha;\downarrow}$ for the spin correlation, $\hat{O}_{j}=\hat{n}_{j,\alpha}-1$ for the charge correlation, and $\hat{O}_{j}=\hat{c}_{j,\alpha;\uparrow}\hat{c}_{j,\alpha;\downarrow}$ for the on-site pairing correlation.
@@
The central result of this Letter is the finding of the DH pairing phase in a finite region of doublon density $n_{d}=\sum_{j,\alpha}\braket{\hat{n}_{j,\alpha;\uparrow}\hat{n}_{j,\alpha;\downarrow}}/(2L)$ and at sufficiently strong nonlocal repulsion $V$, shown in Fig. 1(a). This DH pairing state is characterized by the correlation function $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ with $\hat{\Delta}^{\mathrm{r}}_{j}=\sum_{\alpha,\sigma}(-1)^{\alpha}\left(1-\hat{n}_{j,\bar{\alpha};\bar{\sigma}}\right)\hat{c}^{\dagger}_{j,\bar{\alpha};\sigma}\hat{c}_{j,\alpha;\sigma}\hat{n}_{j,\alpha;\bar{\sigma}}/2$, where $\bar{\alpha}=1$ $(0)$ for $\alpha=0$ $(1)$, and $\bar{\sigma}=\downarrow$ ()($\uparrow$)( ↑ ) for $\sigma=\uparrow$ ()($\downarrow$)( ↓ ). The operator $\hat{\Delta}^{\mathrm{r}\dagger}_{j}$ creates a rung $\eta$-triplet with a doublon and a holon on the rung spin-singlet background, as shown in Fig. 1(b) [91]. Here, the $\eta$-triplet states ($\eta=1$) with $\eta^{z}=-1,0$, and $+1$ correspond to holon-holon (HH), DH, and doublon-doublon (DD) pair states, respectively [86]. Considering ^$\hat{\mathcal{H}}^{(\eta)}_{J_{\perp}}+\hat{\mathcal{H}}_{V_{\perp}}$ on a single rung, we can see that the formation of the rung DH pair in the $\eta$-triplet state is energetically favorable in the presence of the nearest-neighbor interactions. Namely, the energy of the DH $\eta$-triplet state $\varepsilon(\eta=1,\eta^{z}=0)=-V_{\perp}$ is lower than the energy of the HH and DD pair states, $\varepsilon(\eta=1,\eta^{z}=\pm 1)=+V_{\perp}$, and is also lower than the energy of the $\eta$-singlet state $\varepsilon(\eta=0,\eta^{z}=0)=-V_{\perp}+J_{\perp}$. Therefore, it is energetically preferable that such DH pairing correlations develop as the photoexcited carriers relax into quasi-steady lowest-energy conditions. In the Supplemental Material [86], we show that DD and HH pairing correlations are indeed weaker and shorter-ranged.
Figure 2 shows the correlation functions at $n_{d}=0.2$ and $V=0.2t_{\parallel}$ in the isotropic ladder ($t_{\perp}=t_{\parallel}$). The log-log plot in Fig. 2(a) exposes that the DH pairing correlation exhibits power-law decay, while the spin, charge, and on-site pairing correlations decay exponentially, indicating that the DH pairing correlation is dominant. As shown in Fig. 2(b), the DH pairing correlation shows a sign alternation along the chain direction. The mechanism of this striped correlation will be discussed later.

### v2

# Doublon-holon pairing state in photodoped Mott insulators

@@ Abstract
We demonstrate the existence of an unconventional pairing state in photodoped Mott insulators on ladder and quasi-two-dimensional geometries, characterized by quasi-long-range doublon-holon correlations that signal Mott exciton condensation. The doublon-holon pairing exhibits correlations of $d$-wave-like symmetry, reminiscent of superconducting pairing in chemically doped Mott insulators. By constructing the phase diagram, using density matrix renormalization group, we reveal that the doublon-holon pairing state in the photodoped ladder emerges between the spin-singlet, charge-density-wave, and $\eta$-pairing phases. Our study suggests that the interplay of charge, spin, and $\eta$-spin degrees of freedom can give rise to exotic quantum many-body states in photodoped Mott insulators.
@@
Introduction. External field driving is a useful tool for generating and controlling intriguing out-of-equilibrium phenomena in correlated electron systems [1, 2, 3, 4, 5, 6]. Photoinduced phase transitions [7, 8, 9, 10, 11, 12, 13], Floquet engineering [14, 15, 16, 17, 18], and high-harmonic generation [19, 20, 21, 22, 23, 24, 25, 26] have been extensively studied to date. In semiconductors, optical excitation across the band gap generates electron and hole carriers. Similarly, photoexcitation in correlation-driven Mott insulators (MIs) creates two types of carriers, doublons (doubly occupied sites) and holons (empty sites) [see Fig. 1(b)] [27, 28, 6, 29, 30, 31], resulting in a photodoped state distinct from conventional photoexcited semiconductors. In a large-gap MI, the carriers are expected to have long lifetimes, leading to a quasisteady state where doublons and holons coexist [32, 33, 34, 35, 36, 37, 38, 39, 40, 41] and possibly bind in excitons [42, 43, 36, 37, 38, 44], which are distinct from excitons in photodoped semiconductors [45, 46].
@@
Figure 1: (a) Phase diagram of the photodoped ladder as a function of the doublon density $n_{d}$ and the intersite Coulomb interaction $V$ with the blue, red, green, and purple points representing the predominance of the spin, $\eta$-pairing, CDW, and doublon-holon (DH) pairing correlations, respectively. At the orange point ($n_{d}=0.5$ and $V=0.2t_{\parallel}$), the $\eta$-pairing and CDW states are degenerate. Parameters: $t_{\perp}=t_{\parallel}$, $J_{\perp}=J_{\parallel}=0.4t_{\parallel}$, and $V=V_{\parallel}=V_{\perp}$. (b) DH pairing state in the photodoped ladder with doublons and holons generated by light. (c) Spin interaction that acts between up and down spins, and $\eta$-spin interaction that acts between a doublon and a holon. (d) Operator that maps a spin singlet on nearest-neighbor sites to a DH $\eta$ triplet.
@@
Photodoping can be a trigger for creating exotic many-body states that are inaccessible in equilibrium, such as the $\eta$-pairing state, which is characterized by long-range correlation of on-site electronic pairs [47, 48, 49, 50, 51, 52, 53, 54, 55, 56]. The photodoped states have been theoretically investigated in detail in the one-dimensional (1D) Hubbard chain [40, 41] and in high-dimensional Hubbard systems [57, 58, 59]. In the former case, the phase diagram comprises the spin-density-wave (SDW) phase at low doublon density $n_{d}$ and the charge-density-wave (CDW) and $\eta$-pairing phases from intermediate to high $n_{d}$ [40, 41]. Although even richer physical properties are anticipated in the two-dimensional (2D) systems because of the increased geometric degrees of freedom and the lack of spin-charge-$\eta$ separation [41], these properties of photodoped 2D MIs are yet to be understood.
@@
Ladder geometry can serve as a bridge between 1D and 2D systems and is more manageable for numerical approaches. For example, density matrix renormalization group (DMRG) calculations, which enable us to obtain numerically precise low-dimensional wave functions, have shown that chemically hole-doped ladders exhibit pair correlations corresponding to $d$-wave superconductivity (SC) in a 2D square lattice [60, 61, 62, 63, 64, 65]. These studies are an important stepping stone to research on $d$-wave SC in the Hubbard and $t$-$J$ models on multileg cylinders, mimicking the 2D square lattice of cuprate superconductors [66, 67, 68, 69, 70, 71]. Furthermore, the pairing mechanism in ladder and bilayer systems [72, 73, 74, 75, 76, 77, 78, 79, 80, 81, 82, 83] has attracted significant attention associated with the recent discovery of high-temperature SC in bilayer nickelates [84, 85, 86, 87]. Because a strong interchain spin coupling favors the formation of spin-singlet rungs, hole doping to the ladder system leads to the formation of interchain hole-hole pairs to minimize the disruption of spin-singlet bonds [72]. Realization of such a pairing mechanism has also been recently proposed with cold atoms in optical lattices [88]. From these considerations, we arrive at the crucial question: what are the equivalent correlation-driven paired states in photodoped MIs?
@@
In this Letter, we show that photodoped MIs in a quasistationary situation exhibit a pairing state composed of doublons and holons (Fig. 1). In the ladder geometry, this pairing state is characterized by the quasi-long-range correlation of doublon-holon (DH) pairs (Mott excitons), suggesting exciton condensation. Due to strong correlations, the state is distinct from excitonic condensates in semiconductors. Namely, we observe opposite phases for the rung and chain pairing, analogous to signatures of superconducting $d$-wave pairing in chemically doped ladders. These findings are further corroborated by results on a multileg cylinder mimicking the 2D square lattice.
@@
Model. To model MIs in the presence of nearest-neighbor interactions, we start with the standard extended Hubbard Hamiltonian
$\hat{\mathcal{H}}_{\mathrm{Hub}}=\sum_{\langle\bm{ij}\rangle}\hat{T}_{\bm{ij}}+U\sum_{\bm{j}}\hat{n}^{\prime}_{\bm{j};\uparrow}\hat{n}^{\prime}_{\bm{j};\downarrow}+\sum_{\langle\bm{ij}\rangle}V_{\bm{ij}}\hat{n}^{\prime}_{\bm{i}}\hat{n}^{\prime}_{\bm{j}},$ (1)
@@
where $\hat{T}_{\bm{ij}}=-t_{\bm{ij}}\sum_{\sigma}\hat{c}^{\dagger}_{\bm{i};\sigma}\hat{c}_{\bm{j};\sigma}+{\rm H.c.}$ is the hopping term. $\hat{c}^{\dagger}_{\bm{j};\sigma}$ ($\hat{c}_{\bm{j};\sigma}$) is the fermion creation (annihilation) operator at site $\bm{j}$ for spin $\sigma=\uparrow,$ $\downarrow$. $\hat{n}_{\bm{j};\sigma}=\hat{c}^{\dagger}_{\bm{j};\sigma}\hat{c}_{\bm{j};\sigma}$, $\hat{n}^{\prime}_{\bm{j};\sigma}=\hat{n}_{\bm{j};\sigma}-1/2$, and $\hat{n}^{\prime}_{\bm{j}}=\hat{n}^{\prime}_{\bm{j};\uparrow}+\hat{n}^{\prime}_{\bm{j};\downarrow}$. Summation $\braket{\bm{i}\bm{j}}$ runs over nearest-neighbor sites. In the two-leg ladder, $\bm{j}=(j,\alpha)$ denotes site at position $j$ on chain $\alpha$ [=0=0 (lower), $1$ (upper)], and we distinguish the hopping integrals $t_{\bm{ij}}=t_{\parallel},t_{\perp}$ and the nearest-neighbor Coulomb interaction $V_{\bm{ij}}=V_{\parallel},V_{\perp}$ along the chain and rung directions, respectively. In the 2D square lattice, $t_{\bm{ij}}=t_{\rm h}$ and $V_{\bm{ij}}=V$. $U$ is the on-site Coulomb repulsion. The lattice constant $a$ is set to one, and we use $t_{\parallel}$ in the two-leg ladder ($t_{\rm h}$ in the square lattice) as the unit of energy unless otherwise specified.
In photodoped MIs, doublons and holons are generated by an optical excitation, as shown in Fig. 1(b). Their recombination rate $\Gamma_{\rm R}\propto\exp(-U/\epsilon_{0})$ is exponentially suppressed with the number of excitations that are emitted in the recombination process [32, 33, 34, 35, 36, 37, 39, 38]. For low densities of doublons and holons, the recombination predominantly happens via emission of magnetic processes, $\epsilon_{0}$ being their characteristic energy [44]. For higher densities, scattering with other doublons and holons ($\epsilon_{0}\sim t_{\bm{ij}}$) can also contribute [33, 34]. Thus, when $U\gg\epsilon_{0}$, the system can be on the timescales that are parametrically long in $U$ considered to be in a pseudoequilibrium state where doublons and holons are conserved [40, 41]. To efficiently consider the case $U\gg\epsilon_{0}$, we introduce an effective model derived by the Schrieffer-Wolff transformation that conserves the number of doublons and holons [36, 37, 40, 89] with the dominant (lowest order) terms:
$\hat{\mathcal{H}}$= $-\sum_{\langle\bm{ij}\rangle,\sigma}t_{\bm{ij}}\left(\hat{d}^{\dagger}_{\bm{i};\sigma}\hat{d}_{\bm{j};\sigma}-\hat{h}^{\dagger}_{\bm{i};\sigma}\hat{h}_{\bm{j};\sigma}+{\rm H.c.}\right)$
$+\sum_{\langle\bm{ij}\rangle}J_{\bm{i}\bm{j}}\left(\hat{\bm{s}}_{\bm{i}}\cdot\hat{\bm{s}}_{\bm{j}}-\frac{1}{4}\delta_{1,\hat{n}_{\bm{i}}}\delta_{1,\hat{n}_{\bm{j}}}\right)$ (2)
$-\sum_{\langle\bm{ij}\rangle}\left[J_{\bm{i}\bm{j}}\left(\hat{\bm{\eta}}_{\bm{i}}\cdot\hat{\bm{\eta}}_{\bm{j}}-\frac{1}{4}\bar{\delta}_{1,\hat{n}_{\bm{i}}}\bar{\delta}_{1,\hat{n}_{\bm{j}}}\right)-4V_{\bm{i}\bm{j}}\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}\right].$
The first term denotes the hopping of doublons and holons, written explicitly with doublon $\hat{d}^{\dagger}_{\bm{j};\sigma}=\hat{c}^{\dagger}_{\bm{j};\sigma}\hat{n}_{\bm{j};\bar{\sigma}}$ and holon $\hat{h}^{\dagger}_{\bm{j};\sigma}=\hat{c}_{\bm{j};\sigma}(1-\hat{n}_{\bm{j};\bar{\sigma}})$ creation operators [36], where $\bar{\sigma}=\downarrow$ $(\uparrow)$ for $\sigma=\uparrow$ $(\downarrow)$. The second term describes the SU(2)(2) symmetric spin-exchange interaction, which is also considered in the traditional $t$-$J$ model. The spin operator is defined as $\hat{\bm{s}}_{\bm{j}}=\sum_{\sigma,\sigma^{\prime}}\hat{c}^{\dagger}_{\bm{j};\sigma}\bm{\sigma}_{\sigma\sigma^{\prime}}\hat{c}_{\bm{j};\sigma^{\prime}}/2$, where $\bm{\sigma}$ is the vector of Pauli matrices. For the hopping $t_{\bm{i}\bm{j}}$, the coupling constant is given by $J_{\bm{i}\bm{j}}=4t^{2}_{\bm{i}\bm{j}}/U>0$ (at $V_{\bm{i}\bm{j}}=0$). Similarly, the third term represents the interaction between doublons and holons [40, 89]. In the absence of $V_{\bm{i}\bm{j}}$, this term reveals the SU(2)(2) symmetry of the Hubbard model, represented by the $\eta$ spin. The $\eta$-spin operator $\hat{\bm{\eta}}_{\bm{j}}$ is given by $\hat{\eta}^{+}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}^{\dagger}_{\bm{j};\downarrow}\hat{c}^{\dagger}_{\bm{j};\uparrow}$, $\hat{\eta}^{-}_{\bm{j}}=(-1)^{j+\alpha}\hat{c}_{\bm{j};\uparrow}\hat{c}_{\bm{j};\downarrow}$, and $\hat{\eta}^{z}_{\bm{j}}=\left(\hat{n}_{\bm{j}}-1\right)/2$ [90]. In Eq. (Doublon-holon pairing state in photodoped Mott insulators), $\bar{\delta}_{1,\hat{n}_{\bm{j}}}=1-\delta_{1,\hat{n}_{\bm{j}}}$, and the $\eta$ term acts only on empty ($\circ$) and doubly occupied ($\uparrow\downarrow$) sites. $\hat{\eta}^{+}_{\bm{i}}\hat{\eta}^{-}_{\bm{j}}$ swaps the positions of neighboring doublon and holon, while $\hat{\eta}^{z}_{\bm{i}}\hat{\eta}^{z}_{\bm{j}}$ measures the energy according to the occupancy status of sites $\bm{i}$ and $\bm{j}$ [91]. The $\eta$-spin interaction in Eq. (Doublon-holon pairing state in photodoped Mott insulators) is ferromagnetic, which preferably forms an $\eta$-spin triplet state at two sites. This $\eta$-spin interaction is activated by photodoping, while it is negligible in the traditional $t$-$J$ model for chemically doped systems. As is clear from Eq. (Doublon-holon pairing state in photodoped Mott insulators), $V_{\bm{i}\bm{j}}$ manifestly breaks the $\eta$-SU(2)(2) symmetry and leads to an anisotropic XXZ-type Hamiltonian for the $\eta$ sector. Overall, $\hat{\mathcal{H}}$ is expected to describe metastable phases in photodoped MIs with a similar level of rigor as the well-known $t$-$J$ model can describe chemically doped materials.
@@
For $U\gg t_{\bm{ij}},J_{\bm{ij}}$, implying slow recombination of doublons and holons, photoexcited carriers have time to relax to the bottom of the upper band. Therefore, we can transiently describe the photodoped state as the lowest-energy state of $\hat{\mathcal{H}}$ at a fixed doublon density $n_{d}$ that corresponds to the excitation protocol. In the two-leg ladder of length $L$, the doublon density is given by $n_{d}=\sum_{j,\alpha}\braket{\hat{n}_{j,\alpha;\uparrow}\hat{n}_{j,\alpha;\downarrow}}/(2L)$. The lowest-energy state, representing the transient photodoped state, is obtained with the DMRG method [92, 93, 94] within a sector with a fixed number of doublons [95]. In the two-leg ladder, the total number of sites is $N_{\rm s}=2L$. The numbers of up- and down-spin fermions are set to $N_{\uparrow}=N_{\downarrow}=N_{\rm s}/2$ (i.e., half filling without spin polarization), while the state is configured as a function of the doublon number. The given doublon number $N_{d}$ determines the doublon density as $n_{d}=N_{d}/N_{\rm s}$. We employ open boundary conditions along the chain ($x$) direction. In the two-leg ladder, we set the chain length to $L=160$ unless otherwise specified. The bond dimension is set to $m=10000$ with the maximum truncation error typically being on the order of $1.0\times 10^{-7}$. For the 2D square lattice, we employ an $8\times 4$ site cylinder with periodic boundary conditions along the $y$ direction, where the bond dimension and the maximum truncation error are $m=15000$ and $1.0\times 10^{-5}$, respectively.
To identify nonequilibrium phases in photodoped states, we analyze correlation functions $C(r)=\braket{\hat{O}^{\dagger}_{j_{0}+r}\hat{O}_{j_{0}}}$ between operators $\hat{O}_{j}$ separated by $r$ sites. The reference site is fixed at $j_{0}=L/4+1$ with open boundary conditions. We present the site-averaged correlation functions in the Supplemental Material [91] and confirm that our conclusion does not rely on the definition of the correlation function. The dominant phase is determined by the correlation function that exhibits the slowest decay with distance $r$.
@@
Numerical results. The central result of this Letter is the finding of a DH pairing phase, characterized with the slowest power-law decay in the correlation functions $\braket{\hat{\Delta}^{\mathrm{r}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$ and $\braket{\hat{\Delta}^{\mathrm{c}\dagger}_{j_{0}+r}\hat{\Delta}^{\mathrm{r}}_{j_{0}}}$, defined via the operators
$\hat{\Delta}^{\mathrm{\mathrm{r}\dagger}}_{j}$ $=\frac{1}{2}\sum_{\alpha=0,1}\sum_{\sigma}(-1)^{\alpha}\hat{d}^{\dagger}_{j,\alpha;\sigma}\hat{h}^{\dagger}_{j,\bar{\alpha};\sigma},$ (3)
$\hat{\Delta}^{\mathrm{\mathrm{c}\dagger}}_{j}$ $=\frac{1}{2}\sum_{\beta=0,1}\sum_{\sigma}(-1)^{\beta}\hat{d}^{\dagger}_{j+\beta,0;\sigma}\hat{h}^{\dagger}_{j+\bar{\beta},0;\sigma},$ (4)
where $\bar{\alpha}=1$ $(0)$ for $\alpha=0$ $(1)$. These operators act on nearest-neighboring sites that are in the spin-singlet state and transfer it to the state with a doublon and a holon in the $\eta$-triplet state, as shown in Fig. 1(d), along the rung (r) and chain (c) directions, respectively. The definitions of the pair operators given above correspond to the formation of a local component of a Mott exciton, energetically favored by the interactions between nearest-neighbor sites (see Supplemental Material [91] for details).